It is known that the average price for a cappuccino in a restaurant is R15.40 with a standard deviation of R1.20. What is the probability that the sampling error for the mean is less than 0.20 for a sample of 31 restaurants?



1
Expert's answer
2021-10-21T12:04:11-0400
"SE=z\times\dfrac{\sigma}{\sqrt{n}}=>z=\dfrac{SE(\sqrt{n})}{\sigma}"

"SE<0.2=>z=\dfrac{0.2(\sqrt{31})}{1.2}"

"P(SE<0.2)=P(Z<\dfrac{0.2(\sqrt{31})}{1.2})"

"\approx P(Z<0.92796)\approx0.823286"

The probability that the sampling error for the mean is less than 0.20 for a sample of 31 restaurants is "0.823286."